Fermi National Laboratory
 

Accelerator Update for April 9, 2007 - April 11, 2007

 

Monday April 9

The day shift began with Operations monitoring the Tevatron (TeV) store 5342 (132.0E30) and the Recycler stash (30.6E10), with the Antiproton Source (Pbar) stacking (18.9E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE power supplies running, but no beam, with NuMI taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

Cryo system technicians reported at 8:31 AM that the TeV sector A2 wet engine was failing. They can’t estimate how long the engine will last, but the repair would take eight hours.

Operations established beam to the Meson MTest experiment around 8:40 AM.

Operations turned off the MiniBooNE power supplies at 8:52 AM and made the area ready for an access.

At 11:01 AM, the Booster chopper tripped off on an interlock fault. Operators investigated and discovered that there was no power to the power supply chassis. Experts began their repairs.

MiniBooNE experts began a horn pulse test around 1:40 PM.

MiniBooNE experts reported at 2:30 PM that they had run the horn for about 20 minutes and everything looked good. They will operate it again later.

Operations terminated store 5342 at 3 PM.

At 3:12 PM, NuMI went down due to a target fan differential pressure problem. Experts began their repairs.

Operations made D0 and CDF ready for access at 3:17 PM.

Cryo system experts began work on the TeV sector A2 wet engine at 3:31 PM.

Operations prepared the TeV for access at 3:56 PM.

Operations reported at 4:07 PM that the Recycler had stashed antiprotons three times during the day shift.

Operations established beam to NuMI at 6:23 PM.

CDF came out of access at 6:37 PM.

MTest experimenters reported at 7:28 PM that they were through with the beam for the night.

TeV and D0 were out of access by 7:58 PM.

A MI expert reported at 9:32 PM that he had completed a MI-8 collimator study.

At 9:47 PM, cryo system techs reported seeing a vacuum problem at TeV sector D1 that started around 5:50 PM.

Operations prepared the TeV for access at 10:42 PM so that vacuum techs could investigate problems at sectors C4 and D1.

At 11:54 PM, a Linac RF station (LRF2) tripped off for the first of two times in five minutes. It reset okay.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Tuesday April 10

The midnight shift began with Operations monitoring the stash (202.1E10), with Pbar stacking (39.8E10), and with NuMI taking beam.

At 12:35 AM, LRF2 tripped off. It reset.

At 1:48 AM, Operations reported a problem with a TeV C0 vacuum microswitch status. A vacuum tech was called in to access C0 and investigate.

At 2:34 AM, LRF3 tripped off. It reset.

At 6:04 AM, LRF2 tripped off. It reset.

Operations established beam to MTest at 6:23 AM.

A vacuum tech reported at 7:51 AM that he found a loose readback connection on the TeV C0 beam valve.

Operations reported at 8:01 AM that the Recycler had stashed three times during the midnight shift.

Operations reported that the TeV access ended at 9:21 AM and operators began the TeV turn on.

Operations put the TeV through a dry squeeze at 11:08 AM.

MiniBooNE reported at 11:39 AM that they had completed a thermal scan of their horn and found one hot bolt. They made arrangements for a repair.

Operations put the TeV through a 4x0 wet squeeze at 11:54 AM.

Operations put the TeV into shot setup at 12:22 PM.

At 12:45 PM, LRF4 tripped off. Beam resumed at 1:07 PM, but then LRF4 tripped off a minute later. Beam resumed at 1:23 PM. Linac experts continued to investigate.

Operations reported at 2:10 PM that they had received permission to send beam to the MiniBooNE experiment.

At 2:25 PM, BRF6 tripped off. It reset.

Operations established store 5345 at 2:30 PM with an initial luminosity of 206.93E30.

MiniBooNE experts reported that they had tightened the loose bolt on their horn. Operations began sending beam to MiniBooNE at 3:10 PM.

Pbar experts began a study period at 3:25 PM.

At 3:32 PM, LRF4 tripped off. Beam resumed eight minutes later.

At 3:46 PM, Operations reported that an Accumulator kicker (A:IKIK) wasn’t pulsing. Operations contacted an expert.

Operations reported at 3:59 PM that the Recycler had stashed two times during the day shift.

At 4:06 PM, BRF6 tripped off. It reset.

At 4:18 PM, MI aborted its beam due to bad beam positions.

A MI expert completed tuning on the MI-8 line at 5 PM.

The MTest experiment reported at 6:11 PM that they were done with the beam. Their experiment had finished its run. Operators disabled the beam line and dropped the MT6 interlocks.

Experts working on A:IKIK reported at 8:48 PM that they had reterminated a PFN cable. The Accumulator was ready to take beam, but they have more work to do at a later time.

At 11 PM, LRF3 tripped off. It reset.

Operations reported at midnight that the Recycler had stashed once during the evening shift.

Wednesday April 11

The midnight shift began with Operations monitoring store 5345 (82.42E30) and the stash (96.3E10), with Pbar stacking (42.5E10), and with NuMI and MiniBooNE taking beam.

At 1:22 AM, BRF6 tripped off for the first of two times in 68 minutes. It reset.

At 4:56 AM, store 5345 aborted due to a ComEd power glitch. The TeV quenched at B11. Many devices tripped off. The stack remained intact and only part of the stash was lost. Operators began the TeV recovery.

Operations began the TeV turn on at 7:46 AM.

Operations reported at 8 AM that the Recycler had stashed two times during the midnight shift.

The Plans for Wednesday

The plans for today are to establish a store. Linac experts need three to four hours to repair and tune LRF2 and the Run Coordinator will allow them the time. The Run Coordinator might also allow Pbar, MI, and Booster work, but at the end of the 9 o’clock meeting that hadn’t been decided. NuMI needs eight hours to flush their kicker’s cooling system; they began the flush this morning.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/11/2007    email Fermilab